Let A, B, and C be three of the most popular television shows of all time. The total number of episodes of these three shows is 581. There are 26 more episodes of A than C, and the difference between the number of episodes of B and C is 12. Find the number of episodes of each show.
Answer by ikleyn(52824)

.

Let x = the number of C.


Then the number of A is x+26

and the number of B is x+12, according to the condition.



You have then this equation for the total


    x + (x+26) + (x+12) = 581


which gives


    3x = 581 - 26 - 12 = 543,  x = 543/3 = 181.


ANSWER.  C = 181;  A = 181 + 26 = 207  and  B = 181 + 12 = 193.

Solved.


Very simple problem.